# Structure and Independence in Hyperbolic Uniform Disk Graphs

### v1

@@ 3.2 Balanced separators
Corollary 3.1 already yields a separator that can be covered with few cliques, if the angle $\varphi$ is not too small. It remains to provide the point $p$ together with the line $m$ through $p$ such that the separator is balanced and $\varphi$ is not too small. For this, let $V$ be a set of $n$ points (the vertices of $G$). We call a point $p\in\mathbb{H}^{2}$ a centerpoint of $V$ if every line through $p$ divides $V$ into two subsets that both have size at least $\frac{n}{3}$. The following lemma provides such a centerpoint (which follows directly from the existence of a Euclidean centerpoint [centerpoint]). It was also observed in [Kisfaludi-Bak21], but we provide a proof for completeness.
{lemma}
[See also [Kisfaludi-Bak21, proof of Lemma 4]] A centerpoint of a set of $n$ points in the hyperbolic plane exists and can be found in $\mathcal{O}(n)$ time.
Proof.
We do this by first converting the points to the Beltrami-Klein model of the hyperbolic plane. Here, any hyperbolic line is also a Euclidean line, which means a Euclidean centerpoint is also a hyperbolic centerpoint. Thus, we can simply find the Euclidean centerpoint (which takes $\mathcal{O}(n)$ time [centerpoint]) and then convert it back to our original model of the hyperbolic plane. ∎
Now, consider the lines $\ell_{1},\dots,\ell_{n}$, where $\ell_{i}$ goes through $p$ and the vertex $v_{i}\in V$. Let without loss of generality $\ell_{1}$ and $\ell_{2}$ be the two consecutive lines with maximum angle between them. As there are $2n$ angles between consecutive lines covering the full $2\pi$ angle, the angle between $\ell_{1}$ and $\ell_{2}$ is at least $\pi/n$. Set $m$ to be the angular bisector of $\ell_{1}$ and $\ell_{2}$. It follows that $\frac{1}{\varphi}\in\mathcal{O}(n)$ and thus Corollary 3.1 yields a separator that can be covered by $\mathcal{O}(\log n\cdot(1+\frac{1}{r}))$ cliques. Moreover, as $m$ goes through the centerpoint $p$, this separator is balanced. Clearly, the line $m$ and thus the separator can be computed in $\mathcal{O}(n\log n)$ time. This concludes the proof of Theorem 1.
See 1
@@ 4 Outerplanarity of hyperbolic Delaunay complexes
Let $S$ be a set of $n$ sites in $\mathbb{H}^{2}$ with pairwise distance at least $2r$. Note that $S$ interpreted as hyperbolic uniform disk graph with radius $r$ forms an independent set. Here we prove Theorem 2, giving an upper bound on the outerplanarity of the Delaunay complex $\mathcal{D}(S)$. We use two different types of arguments. The first argument is, roughly speaking, based on the observation that any inner vertex of $\mathcal{D}(S)$ requires many other sites around it to shield it from being an outer vertex. This is similar to the observation in the introduction (recall Figure 1) that for high radius $r$, large stars can be realized. Consequentially, this first type of argument only works in case $r$ is a sufficiently large constant. The second type of argument considers layers of bounded Voronoi cells (which correspond to inner vertices of $\mathcal{D}(S)$) around one fixed center cell. As the union of these layers is bounded by a polygon with a linear number of vertices, its area is linear. In contrast to that, we will see that the area of the layers grows exponentially with the layer (with the base depending on $r$), showing that there cannot be too many layers.
Before we start with the proof, we make one more observation. If the sites $S$ are in general position, i.e., no four sites lie on the same circle, then the Voronoi diagram is a $3$-regular graph and the Delaunay complex is internally triangulated, i.e., each inner face is a triangle. Otherwise, the sites $S$ can be slightly perturbed to give a set $S^{\prime}$, such that $\mathcal{D}(S^{\prime})$ is obtained from $\mathcal{D}(S)$ by triangulating each inner face. Moreover $\mathcal{V}(S^{\prime})$ is obtained from $\mathcal{V}(S)$ by splitting each vertex of degree more than $3$ into a binary tree. Triangulating $\mathcal{D}(S)$ only adds edges to it and thereby only increases its outerplanarity. Thus, any upper bound on the outerplanarity of $\mathcal{D}(S^{\prime})$ also holds for $\mathcal{D}(S)$. For this section, we assume without loss of generality that $S$ is in general position, i.e., $\mathcal{D}(S)$ is internally triangulated.
@@ 4.1 Large-radius case
Therefore, the angle between $sa$ and $sb$ is at most $2\alpha\leq 2\pi e^{-r}$. It follows that $s$ has at least $\frac{2\pi}{2\pi e^{-r}}=e^{r}$ neighbors. ∎
@@
It follows that removing all outer vertices yields a plane graph with less than $n\cdot 6/d$ vertices in which every inner vertex again has degree at least $d$. Thus, repeatedly removing all outer vertices $k$ times yields a graph with less than $n\cdot(6/d)^{k}$ vertices. Hence $G$ is reduced to at most one vertex in less than $\log_{d/6}n$ rounds, which concludes the proof. ∎
For $r\geq 1.8>\log 6$, these two lemmas already yield the claim of Theorem 2; also see the formal proof in Section 4.3, where we combine the large- and small-radius case.
@@ 4.2 Small-radius case
As mentioned above, we use an argument based on area for the case that the radius is a small constant or even decreasing with $n$. Before we start, we give a simple area-based argument for why $r\geq\log n$ implies that the Delaunay complex has no inner vertices. Although our argument for small radii is more involved, this gives a good intuition how the area behaves in the hyperbolic plane and how this can help us in proving outerplanarity. Let $s\in S$ be a site. If $s$ is an inner vertex of $\mathcal{D}(S)$, then its Voronoi cell is bounded and its boundary is a polygon of at most $n-1$ vertices. Thus, the area of this Voronoi cell is less than $(n-3)\pi$. Simultaneously, the disk of radius $r$ around $s$ is included in the Voronoi cell of $s$ as all sites have pairwise distance at least $2r$. The area of this disk is $4\pi\sinh^{2}(\frac{r}{2})$ which is larger than $e^{r}$ for sufficiently large $r$. From this it follows that bounded cells can only exist when $r<\log n$.
To make an argument that works for smaller $r$, let $s\in S$ be an arbitrary but fixed vertex of the Delaunay complex $\mathcal{D}(S)$. We partition the vertices of $\mathcal{D}(S)$ into layers by hop distance from $s$ in $\mathcal{D}(S)$, i.e., $V_{\ell}$ is the set of vertices with distance $\ell$ from $s$. Let $L$ be the largest integer such that for all $\ell\leq L$ the layer $V_{\ell}$ contains only inner vertices. Note that our goal is to prove an upper bound on $L$ as this bounds the distance of $s$ to the outer face.
Before formalizing the above proof idea, we show two simple results about hyperbolic triangle areas, both based on the fact that a right-angled triangle with short sides of length $a$ and $b$ has area $2\arctan\left(\tanh\frac{a}{2}\cdot\tanh\frac{b}{2}\right)$. We start with an upper bound on the area of a triangle where one side has a specific length. This will serve as an upper bound for the triangles in Figure 6(b).
@@
$2\arctan\left(\tanh\frac{a_{1}}{2}\tanh\frac{b}{2}\right)+2\arctan\left(\tanh\frac{a_{2}}{2}\tanh\frac{b}{2}\right)$ $\leq 2\arctan\left(\tanh\frac{a_{1}}{2}\right)+2\arctan\left(\tanh\frac{a_{2}}{2}\right)$
@@
Proof.
The area is given by the function $f_{b}(a):=2\arctan\left(\tanh\frac{a}{2}\cdot\tanh\frac{b}{2}\right)$, which is concave, as $f_{b}^{\prime\prime}(a)=-\frac{2\sinh a\sinh(2b)}{(2+2\cosh a\cosh b)^{2}}\leq 0$ for any $a,b\in\mathbb{R}$. Because additionally $f_{b}(0)=0$, we can say that $f_{b}(a)\geq a\cdot f_{b}(1.8)/1.8$ for $a\in[0,1.8]$. What remains is to bound $g(b):=f_{b}(1.8)$. This function $g(b)$ is concave for analogous reasons as before, and again $g(0)=0$, so we can say that $g(b)\geq b\cdot g(\pi)/\pi\geq 0.37b$ for $b\in[0,\pi]$. Additionally, $g(b)\geq 0.37\pi$ for $b\geq\pi$. Thus, $f_{b}(a)\geq a/5\cdot\min\{b,\pi\}$ for $a\in[0,1.8]$. ∎
@@
Observe that this lower bound is the same as the upper bound except for the factor of $r/10$. Thus, we obtain $\operatorname{\mathrm{Area}}(P_{\ell})-\operatorname{\mathrm{Area}}(P_{\ell-1})\geq\frac{r}{10}\operatorname{\mathrm{Area}}(P_{\ell})$. Rearranging yields the claimed bound. ∎
@@ 4.3 Combining the two
First note that $\log\frac{10}{10-r}\geq\frac{r}{10}$ for $r<10$ and thus it remains to show that $\frac{\operatorname{\mathrm{Area}}(P_{L})}{\operatorname{\mathrm{Area}}(P_{1})}$ is polynomial in $n$. For this observe that $\operatorname{\mathrm{Area}}(P_{L})\leq 2\pi n$ as it is a polygon with less than $2n$ vertices. Moreover, $P_{1}$ at least includes the Voronoi cell of $s$, which includes a disk of radius $r$ as sites have pairwise distance at least $2r$. Thus, $\operatorname{\mathrm{Area}}(P_{1})\geq 4\pi\sinh^{2}(r/2)\in\Omega(1/n^{2})$ as $r\geq 1/n$ and $\sinh x$ behaves like $x$ for $x$ close to 00. It follows that $L\in\mathcal{O}\big{(}\frac{\log n}{r}\big{)}$, which concludes the proof. ∎
@@ 5 Exact algorithm for Independent Set
In the following, we give an exact algorithm for computing an independent set of a given size in a hyperbolic uniform disk graph $G$ of radius $r$ that is given via its geometric realization. Let $S$ be an independent set of $G$. As there are no edges between vertices in $S$, they have pairwise distance at least $2r$. Thus, Theorem 2 gives us an upper bound on the outerplanarity of the Delaunay complex $\mathcal{D}(S)$, which implies that $\mathcal{D}(S)$ has small treewidth and thereby small balanced separators. In a nutshell, our algorithm first lists all relevant separators of the Delaunay complexes of all possible independent sets. We then use a dynamic program to combine these candidate separators into a hierarchy of separators, maximizing the number of vertices of the corresponding Delaunay complex and thereby the size of the independent set $S$.
@@ 5.1 Sphere cut decompositions
A branch decomposition [Robertson_Seymour_1991] of a graph $G$ is an unrooted binary tree $T$, i.e., each non-leaf node has degree exactly 3, and there is a bijection between the leaves of $T$ and the edges of $G$. Observe that removing an edge $e_{T}\in E(T)$ separates $T$ into two subtrees and thereby separates its leaves and thus the edges of $G$ in two subsets. Let $E_{1}\subset E(G)$ and $E_{2}\subset E(G)$ be these two edge sets and let $G[E_{1}]$ and $G[E_{2}]$ be their induced subgraphs. The vertices shared by $G[E_{1}]$ and $G[E_{2}]$, i.e., vertices that appear in edges of $E_{1}$ and of $E_{2}$, are called the middle set of $e_{T}$. The width of the branch decomposition $T$ is the size of the largest middle set over all edges of $T$. The branch-width of $G$ is the minimum width of all branch decompositions of $G$. Robertson and Seymour [Robertson_Seymour_1991] showed that the branch-width of a graph is within a constant factor of its treewidth.
@@
If $G$ is a plane graph, results by Seymour and Thomas [DBLP:journals/combinatorica/SeymourT94] imply that we can wish for some additional structure without increasing the width of the branch decomposition [Dorn_Penninkx_Bodlaender_Fomin_2010, MarxP22]. Specifically, a sphere cut decomposition of a plane graph $G$ is a branch decomposition $T$ such that every tree edge $e_{T}\in E(T)$ is associated with a so-called noose. The noose $O$ of $e_{T}$ is a simple closed curve that intersects the vertices of $G$ exactly in the middle set of $e_{T}$ and separates $G[E_{1}]$ from $G[E_{2}]$; see also Figure 8. Moreover, removing the vertices of the middle set from $O$ separates $O$ into noose segments such that each segment lies entirely in the interior of a face of $G$ and every face contains at most one noose segment, except if $E_{1}$ or $E_{2}$ contains just a bridge of $G$. In the latter case, the noose consists of two noose segments through the face incident to the bridge. It should be mentioned that this differs from the original definition by Dorn, Penninkx, Bodlaender and Fomin [Dorn_Penninkx_Bodlaender_Fomin_2010]. Here, a noose may intersect any face at most once and thus the graph cannot have bridges, as discussed by Marx and Pilipczuk [MarxP22]. With our relaxed definition of nooses however, we get the following theorem.
Theorem 5 ([Dorn_Penninkx_Bodlaender_Fomin_2010, MarxP22]).
@@ 5.2 Geometric nooses of Delaunay complexes
Figure 9: Part (a): Delaunay complex $\mathcal{D}(S)$ (black) together with the corresponding Voronoi diagram (blue). Part (b) visualizes the choice of the ideal point $p_{u}$ for a Voronoi ray, as needed for the normalization of the green noose. Part (c) shows the normalized version of the noose from part (b) alongside two other normalized nooses, each separating a single edge of $\mathcal{D}(S)$. Observe that the depicted normalized nooses include zero, one, or two ideal arcs, respectively. Part (d) shows a parent noose (green) and its child nooses, with interiors colored orange and blue.
@@
If $f=f_{\infty}$ is the outer face, $u$ and $v$ can have multiple incidences to $f_{\infty}$ (in fact $u$ and $v$ could be the same vertex). To make the situation more precise, assume that we traverse the boundary of $f_{\infty}$ such that $f_{\infty}$ lies to the left and let $e_{u}$ be the edge incident to $f_{\infty}$ that precedes the incidence of $u$ where the noose enters $f_{\infty}$; see Figure 9(b). Recall that the dual Voronoi edge $e_{u}^{\star}$ is unbounded and ends in some ideal point $p_{u}$. Let $e_{v}$ and $p_{v}$ be defined analogously for $v$. Then, for the normalized noose segment from $u$ to $v$ in $f_{\infty}$, we use the ray between $u$ and $p_{u}$, the ideal arc from $p_{u}$ to $p_{v}$, and the ray between $p_{v}$ and $v$.
Observe that combining the geometric noose segments as defined above yields a generalized polygon for each noose; see Figure 9(c). Also, note that this definition also works in the special case where the noose goes only through one vertex with two incidences to the outer face, in which case the noose consists of just two rays and an ideal arc.
@@
The previous lemma summarizes the core properties of individual nooses. Next, we describe additional properties of how nooses interact in their parent-child relationships. See also Figure 9(d).
{lemma}
@@
Proof.
To prove this, we consider how the nooses separate the graph. There is an internal node of the sphere cut decomposition that is incident to the three edges associated with $O_{P}$, $O_{L}$, and $O_{R}$ and this internal node separates the edges of $\mathcal{D}(S)$ into three parts $E_{P}$, $E_{L}$, and $E_{R}$ such that in $\mathcal{D}(S)$ the noose $O_{P}$ separates $E_{L}\cup E_{R}$ from $E_{P}$, $O_{L}$ separates $E_{L}$ from $E_{R}\cup E_{P}$, and $O_{R}$ separates $E_{R}$ from $E_{P}\cup E_{L}$.
The faces visited by $O_{L}$ are exactly those faces of $\mathcal{D}(S)$ whose boundary contains edges both from $E_{L}$ and from $E_{R}\cup E_{P}$, and analogous statements hold for $O_{R}$ and $O_{P}$. This means that any noose that separates, e.g., $E_{L}$ from the rest of the graph has to pass through a uniquely determined set of vertices and faces in a uniquely determined cyclic order. For two nooses that visit a shared set of faces, the cyclic order in which these faces and the incident vertices are visited thus has to coincide. We can thus consider a combinatorial representation $O^{\star}_{L}$, $O^{\star}_{R},$ and $O^{\star}_{P}$ of the three nooses as a cyclic order of vertex-face-incidences and it follows that $O^{\star}_{P}$ is the symmetric difference of $O^{\star}_{L}$ and $O^{\star}_{R}$. For a vertex-face-incidence of a combinatorial noose, its normalized geometric realization contains a line segment or ray that is uniquely determined by the vertex and face according to the construction of normalized nooses. This implies that $O_{P}$ is the symmetric difference of $O_{L}$ and $O_{R}$, except for exactly two points that are visited by all three normalized nooses. These correspond either to a vertex of $\mathcal{D}(S)$ at which all three nooses meet or to an (ideal) Voronoi vertex located in a face of $\mathcal{D}(S)$ that is visited by all three nooses. As the nooses are simple closed curves, it directly follows that the interiors of $O_{L}$ and $O_{R}$ are subsets of the interior of $O_{P}$. ∎
@@
{lemma}
Let $S$ be a set of sites pairwise distance at least $2r$, let $\mathcal{D}(S)$ be its Delaunay complex, and let $O$ be a noose of a normalized sphere cut decomposition of $\mathcal{D}(S)$. Then $O$ has distance at least $r$ from any site not visited by $O$.
Proof.
@@
As this holds for any two consecutive sites visited by $O$, it follows that $O$ goes only through Voronoi cells of sites it visits. Let $s\in S$ be a site not visited by $O$. As all other sites have pairwise distance at least $2r$ to $s$, the open disk of radius $r$ around $s$ lies inside the Voronoi cell of $s$. As the noose does not enter the Voronoi cell of $s$, it does not enter this open disk and thus has distance at least $r$ from $s$. ∎

### v2

@@ 5.2 Balanced separators
Corollary 5.5 already yields a separator that can be covered with few cliques, if the angle $\varphi$ is not too small. It remains to provide the point $p$ together with the line $m$ through $p$ such that the separator is balanced and $\varphi$ is not too small. For this, let $V$ be a set of $n$ points (the vertices of $G$). We call a point $p\in\mathbb{H}^{2}$ a centerpoint of $V$ if every line through $p$ divides $V$ into two subsets that both have size at least $\frac{n}{3}$. The following lemma provides such a centerpoint (which follows directly from the existence of a Euclidean centerpoint [26]). It was also observed in [28], but we provide a proof for completeness.
Lemma 5.6 (See also [28, proof of Lemma 4]).
A centerpoint of a set of $n$ points in the hyperbolic plane exists and can be found in $\mathcal{O}(n)$ time.
Proof 5.7.
We do this by first converting the points to the Beltrami-Klein model of the hyperbolic plane. Here, any hyperbolic line is also a Euclidean line (see also Figure 3). This means that a Euclidean centerpoint is also a hyperbolic centerpoint. Thus, we can simply find the Euclidean centerpoint (which takes $\mathcal{O}(n)$ time [26]) and then convert it back to our original model of the hyperbolic plane.
Now, consider the lines $\ell_{1},\dots,\ell_{n}$, where $\ell_{i}$ goes through $p$ and the vertex $v_{i}\in V$. Let without loss of generality $\ell_{1}$ and $\ell_{2}$ be the two consecutive lines with maximum angle between them. As there are $2n$ angles between consecutive lines covering the full $2\pi$ angle, the angle between $\ell_{1}$ and $\ell_{2}$ is at least $\pi/n$. Set $m$ to be the angular bisector of $\ell_{1}$ and $\ell_{2}$. It follows that $\frac{1}{\varphi}\in\mathcal{O}(n)$ and thus Corollary 5.5 yields a separator that can be covered by $\mathcal{O}(\log n\cdot(1+\frac{1}{r}))$ cliques. Moreover, as $m$ goes through the centerpoint $p$, this separator is balanced. Clearly, the line $m$ and thus the separator can be computed in $\mathcal{O}(n\log n)$ time. This concludes the proof of Theorem 1.1.
See 1.1
@@ 6 Outerplanarity of hyperbolic Delaunay complexes
Before we start with the proof, we make an observation. If the sites $S$ are in general position, i.e., no four sites lie on the same circle, then the Voronoi diagram is a $3$-regular graph and the Delaunay complex is internally triangulated, i.e., each inner face is a triangle. Otherwise, we can slightly perturb the sites $S$ to get a set $S^{\prime}$ in general position, such that $\mathcal{D}(S^{\prime})$ is obtained from $\mathcal{D}(S)$ by triangulating each inner face. Moreover $\mathcal{V}(S^{\prime})$ is obtained from $\mathcal{V}(S)$ by splitting each vertex of degree more than $3$ into a binary tree. Triangulating $\mathcal{D}(S)$ only adds edges to it and thereby only increases its outerplanarity. Thus, any upper bound on the outerplanarity of $\mathcal{D}(S^{\prime})$ also holds for $\mathcal{D}(S)$. For this section, we assume without loss of generality that $S$ is in general position, i.e., $\mathcal{D}(S)$ is internally triangulated.
@@ 6.1 Large-radius case
Therefore, the angle between $sa$ and $sb$ is at most $2\alpha\leq 2\pi e^{-r}$. It follows that $s$ has at least $\frac{2\pi}{2\pi e^{-r}}=e^{r}$ neighbors.
@@
It follows that removing all outer vertices yields a plane graph with less than $n\cdot 6/d$ vertices in which every inner vertex again has degree at least $d$. Thus, repeatedly removing all outer vertices $k$ times yields a graph with less than $n\cdot(6/d)^{k}$ vertices. Hence $G$ is reduced to at most one vertex in less than $\log_{d/6}n$ rounds, which concludes the proof.
For $r\geq 1.8>\log 6$, these two lemmas already yield the claim of Theorem 1.4; also see the formal proof in Section 6.3, where we combine the large- and small-radius case.
@@ 6.2 Small-radius case
Before formalizing the proof idea, we show two simple results about hyperbolic triangle areas, both based on the fact that a right-angled triangle with short sides of length $a$ and $b$ has area $2\arctan\left(\tanh\frac{a}{2}\cdot\tanh\frac{b}{2}\right)$. We start with an upper bound on the area of a triangle where one side has a specific length. This will serve as an upper bound for the triangles in Figure 5(b).
@@
$2\arctan\left(\tanh\frac{a_{1}}{2}\tanh\frac{b}{2}\right)$ $+2\arctan\left(\tanh\frac{a_{2}}{2}\tanh\frac{b}{2}\right)$
$\leq 2\arctan\left(\tanh\frac{a_{1}}{2}\right)+2\arctan\left(\tanh\frac{a_{2}}{2}\right)$
@@
Proof 6.8.
The area is given by the function $f_{b}(a):=2\arctan\left(\tanh\frac{a}{2}\cdot\tanh\frac{b}{2}\right)$, which is concave, as $f_{b}^{\prime\prime}(a)=-\frac{2\sinh a\sinh(2b)}{(2+2\cosh a\cosh b)^{2}}\leq 0$ for any $a,b\in\mathbb{R}$. Because additionally $f_{b}(0)=0$, we can say that $f_{b}(a)\geq a\cdot f_{b}(1.8)/1.8$ for $a\in[0,1.8]$. What remains is to bound $g(b):=f_{b}(1.8)$. This function $g(b)$ is concave for analogous reasons as before, and again $g(0)=0$, so we can say that $g(b)\geq b\cdot g(\pi)/\pi\geq 0.37b$ for $b\in[0,\pi]$. Additionally, $g(b)\geq 0.37\pi$ for $b\geq\pi$. Thus, $f_{b}(a)\geq a/5\cdot\min\{b,\pi\}$ for $a\in[0,1.8]$.
@@
Observe that this lower bound is the same as the upper bound except for the factor of $r/10$. Thus, we obtain $\operatorname{\mathrm{Area}}(P_{\ell})-\operatorname{\mathrm{Area}}(P_{\ell-1})\geq\frac{r}{10}\operatorname{\mathrm{Area}}(P_{\ell})$. Rearranging yields the claimed bound.
@@ 6.3 Combining the two
First note that $\log\frac{10}{10-r}\geq\frac{r}{10}$ for $r<10$ and thus it remains to show that $\frac{\operatorname{\mathrm{Area}}(P_{L})}{\operatorname{\mathrm{Area}}(P_{1})}$ is polynomial in $n$. For this observe that $\operatorname{\mathrm{Area}}(P_{L})\leq 2\pi n$ as it is a polygon with less than $2n$ vertices. Moreover, $P_{1}$ at least includes the Voronoi cell of $s$, which includes a disk of radius $r$ as sites have pairwise distance at least $2r$. Thus, $\operatorname{\mathrm{Area}}(P_{1})\geq 4\pi\sinh^{2}(r/2)\in\Omega(1/n^{2})$ as $r\geq 1/n$ and $\sinh x$ behaves like $x$ for $x$ close to 00. It follows that $L\in\mathcal{O}\big{(}\frac{\log n}{r}\big{)}$, which concludes the proof.
@@ 7 Exact algorithm for independent set
In the following, we give an exact algorithm for computing an independent set of a given size in a hyperbolic uniform disk graph $G$ of radius $r$ that is given via its geometric realization. Let $S$ be an independent set of $G$. As there are no edges between vertices in $S$, they have pairwise distance at least $2r$. Thus, Theorem 1.4 gives us an upper bound on the outerplanarity of the Delaunay complex $\mathcal{D}(S)$, which implies that $\mathcal{D}(S)$ has small treewidth and thereby small balanced separators. In a nutshell, our algorithm first lists all relevant separators of the Delaunay complexes of all possible independent sets. We then use a dynamic program to combine these candidate separators into a hierarchy of separators, maximizing the number of vertices of the corresponding Delaunay complex and thereby the size of the independent set $S$.
@@ 7.1 Sphere cut decompositions
A branch decomposition [43] of a graph $G$ is an unrooted binary tree $T$, i.e., each non-leaf node has degree exactly 3, and there is a bijection between the leaves of $T$ and the edges of $G$. Observe that removing an edge $e_{T}\in E(T)$ separates $T$ into two subtrees and thereby separates its leaves and thus the edges of $G$ in two subsets. Let $E_{1}\subset E(G)$ and $E_{2}\subset E(G)$ be these two edge sets and let $G[E_{1}]$ and $G[E_{2}]$ be their induced subgraphs. The vertices shared by $G[E_{1}]$ and $G[E_{2}]$, i.e., vertices that appear in edges of $E_{1}$ and of $E_{2}$, are called the middle set of $e_{T}$. The width of the branch decomposition $T$ is the size of the largest middle set over all edges of $T$. The branch-width of $G$ is the minimum width of all branch decompositions of $G$. Robertson and Seymour [43] showed that the branch-width of a graph is within a constant factor of its treewidth.
@@
If $G$ is a plane graph, results by Seymour and Thomas [44] imply that we can wish for some additional structure without increasing the width of the branch decomposition [17, 37]. Specifically, a sphere cut decomposition of a plane graph $G$ is a branch decomposition $T$ such that every tree edge $e_{T}\in E(T)$ is associated with a so-called noose. The noose $O$ of $e_{T}$ is a simple closed curve that intersects the vertices of $G$ exactly in the middle set of $e_{T}$ and separates $G[E_{1}]$ from $G[E_{2}]$; see also Figure 10. Moreover, removing the vertices of the middle set from $O$ separates $O$ into noose segments such that each segment lies entirely in the interior of a face of $G$ and every face contains at most one noose segment, except if $E_{1}$ or $E_{2}$ contains just a bridge of $G$. In the latter case, the noose consists of two noose segments through the face incident to the bridge. It should be mentioned that this differs from the original definition by Dorn, Penninkx, Bodlaender and Fomin [17]. Here, a noose may intersect any face at most once and thus the graph cannot have bridges, as discussed by Marx and Pilipczuk [37]. With our relaxed definition of nooses however, we get the following theorem.
Theorem 7.1 ([17, 37]).
@@ 7.2 Geometric nooses of Delaunay complexes
Figure 11: Part (a) visualizes the choice of the ideal point $p_{u}$ for a Voronoi ray, as needed for the normalization of the green noose. Part (b) shows the normalized version of the noose from part (a) alongside two other normalized nooses, each separating a single edge of $\mathcal{D}(S)$. Observe that the depicted normalized nooses include zero, one, or two ideal arcs, respectively. Part (c) shows a parent noose (blue) and its child nooses, with interiors colored green and orange.
@@
If $f=f_{\infty}$ is the outer face, $u$ and $v$ can have multiple incidences to $f_{\infty}$ (in fact $u$ and $v$ could be the same vertex). To make the situation more precise, assume that we traverse the boundary of $f_{\infty}$ such that $f_{\infty}$ lies to the left and let $e_{u}$ be the edge incident to $f_{\infty}$ that precedes the incidence of $u$ where the noose enters $f_{\infty}$; see Figure 11(a). Recall that the dual Voronoi edge $e_{u}^{\star}$ is unbounded and ends in some ideal point $p_{u}$. Let $e_{v}$ and $p_{v}$ be defined analogously for $v$. Then, for the normalized noose segment from $u$ to $v$ in $f_{\infty}$, we use the ray between $u$ and $p_{u}$, the ideal arc from $p_{u}$ to $p_{v}$, and the ray between $p_{v}$ and $v$.
Observe that combining the geometric noose segments as defined above yields a generalized polygon for each noose; see Figure 11(b). Also, note that this definition also works in the special case where the noose goes only through one vertex with two incidences to the outer face, in which case the noose consists of just two rays and an ideal arc.
@@
The previous lemma summarizes the core properties of individual nooses. Next, we describe additional properties of how nooses interact in their parent-child relationships. See also Figure 11(c).
Lemma 7.3.
@@
Proof 7.4.
Recall that a noose is a simple closed curve intersecting $G$ precisely at its associated vertex separator. It can be viewed as a sequence of segments, each traversing a face by connecting two incident vertices. For the proof of the lemma, we first take a step back from geometric perspective and define the combinatorial representation of a noose as the (multi-)set of its vertex–face incidences. Note that the combinatorial representation of noose $O$ is already uniquely determined by its interior and exterior edges.
We show that the combinatorial representation of $O_{P}$ is the symmetric difference of the combinatorial representations of $O_{L}$ and $O_{R}$. For this, we partition the edges of $G$ into edges $E_{L}$ in the interior of $O_{L}$, $E_{R}$ in the interior of $O_{R}$ and $E_{\mathrm{ex}}$ in the exterior of $O_{P}$. Then the edges in the interior of $O_{P}$ are $E_{L}\cup E_{R}$. We consider a vertex–face incidence $i=(v,f)$ of one of the child nooses, without loss of generality $O_{L}$. As $O_{L}$ separates $E_{L}$ from $E_{R}\cup E_{\mathrm{ex}}$, next to $i$ on the boundary of $f$ there is an edge $e_{1}\in E_{L}$ and an edge $e_{2}\in E_{R}\cup E_{\mathrm{ex}}$. If $e_{2}\in E_{R}$, then $O_{R}$ also enters $f$ via $v$ and also contains the vertex–face incidence $i$, while $O_{P}$ does not contain $i$. If otherwise $e_{2}\in E_{\mathrm{ex}}$, then $O_{P}$ also visits $f$ through the same vertex incidence, while $O_{R}$ does not. All other cases are analogous and in particular any vertex–face incidence of the parent noose occurs in exactly one of the child nooses. It follows that the vertex–face incidences of $O_{P}$ are the symmetric difference of those of $O_{L}$ and $O_{R}$.
The normalized geometric realization of a noose contains a line segment or ray for every vertex–face incidence. This line segment or ray is uniquely determined by the incidence’s vertex and face according to the construction of normalized nooses. As $O_{P}$, $O_{L}$, and $O_{R}$ are nooses of a normalized sphere cut decomposition, it follows that $E(O_{P})$ is the symmetric difference of $E(O_{L})$ and $E(O_{R})$. [^11] Recall that $O_{P}$ is a generalized polygon and thus $E(O_{P})$ refers to the edges of a plane graph. Thus, $O_{P}$ is the symmetric difference of $O_{L}$ and $O_{R}$, except for exactly two points that are visited by all three normalized nooses. These correspond either to a vertex of $\mathcal{D}(S)$ at which all three nooses meet or to an (ideal) Voronoi vertex located in a face of $\mathcal{D}(S)$ that is visited by all three nooses. As the nooses are simple closed curves, it directly follows that the interiors of $O_{L}$ and $O_{R}$ are subsets of the interior of $O_{P}$.
@@
Lemma 7.5.
Let $S$ be a set of sites with pairwise distance at least $2r$, let $\mathcal{D}(S)$ be its Delaunay complex, and let $O$ be a noose of a normalized sphere cut decomposition of $\mathcal{D}(S)$. Then $O$ has distance at least $r$ from any site not visited by $O$.
Proof 7.6.
@@
As this holds for any two consecutive sites visited by $O$, it follows that $O$ goes only through Voronoi cells of sites it visits. Let $s\in S$ be a site not visited by $O$. As all other sites have pairwise distance at least $2r$ to $s$, the open disk of radius $r$ around $s$ lies inside the Voronoi cell of $s$. As the noose does not enter the Voronoi cell of $s$, it does not enter this open disk and thus has distance at least $r$ from $s$.